# A Priori Uncertainty Quantification of Reacting Turbulence Closure Models using Bayesian Neural Networks

### v1

@@ 3.3 Validation and Interpretation in Physical Space
The predicted total filtered progress variable dissipation rates, $\widetilde{\chi}=\chi_{\widetilde{C}}+\chi_{\rm SFS}$ and associated uncertainties (epistemic and aleatoric) are visualized for two different flames: an n-heptane flame with Karlovitz number $7$ savard2017effects which was not included in the dataset (referred to as $B_{\rm Le}$), and a n-heptane flame with Karlovitz number $256$ which was included in the dataset (referred to as $D_{\rm Le}$). The prediction results and the ground truth data are shown for filter width 4 (Fig. 9) and 16 (Fig. 10).
@@ 3.4 Extrapolation Behavior
$\|\mu_{\rm OOD}-\mathbb{E}_{q_{c}(\mathbf{w}|\theta)}\left[\mathbf{y}|\mathbf{x},\boldsymbol{D_{\rm OOD}}\right]\|_{2}/\sqrt{N},$ (11)
where $N$ is the number of synthetic data points outside of the data distribution, and the normalization $1/\sqrt{N}$ accounts for the different synthetic dataset sizes. Likewise, the normalized $L_{2}$ error for the second moment is computed as
$\|\sigma_{\rm OOD}^{2}-\operatorname{Var}\left(\mathbb{E}_{q_{c}(\mathbf{w}|\theta)}\left[\mathbf{y}|\mathbf{x},\boldsymbol{D_{\rm OOD}}\right]\right)\|_{2}/\sqrt{N}.$ (12)
@@
With the addition of increasing amounts of synthetic data, the model is tasked with representing a larger space and it is natural to expect that the performance should degrade on the original dataset. Figure 12 shows the effect of the amount of synthetic data on the relative $L_{2}$ error on the original dataset. For a given model $M$, the predictive mean is noted as $P_{M}=\mathbb{E}_{q_{M,c}(\mathbf{w}|\theta)}\left[\mathbf{y}|\mathbf{x},\boldsymbol{D}\right]$. The error metric considered is a relative error between the mean predictions of a model trained without synthetic data $P_{\rm ref}$ and a model trained with different amounts of synthetic data $P_{\rm OOD}$, as
@@
4 Conclusion
@@
A strategy for enforcing a specific OOD behavior was proposed using synthetic data generation. It was found that using a uniform in-phase space data generation led to the best performances and that the performance in-distribution did not degrade until the synthetic dataset size approached that of the original dataset.
@@ 5 Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0021110. This work was authored in part by the National Renewable Energy Laboratory (NREL), operated by Alliance for Sustainable Energy, LLC, for the U.S. Department of Energy (DOE) under Contract No. DE-AC36-08GO28308. This work was supported as part of DEGREES, an Energy Earthshot Research Center (EERC) funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences. The research was performed using computational resources sponsored by the Department of Energy’s Office of Energy Efficiency and Renewable Energy and located at the National Renewable Energy Laboratory. The views expressed in the article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes.
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
@@ Appendix B Computation of Uncertainties
In this appendix, the algorithms for the computation of a predictive distribution (Alg. 2), the epistemic uncertainty (Alg. 3) and the aleatoric uncertainty (Alg. 4) are presented. Overall, the BNN allows for fast evaluation of the samples of weights, but appropriate averaging is needed to differentiate between epistemic and aleatoric uncertainties.
@@
Computation of the epistemic uncertainty relies on the parameterization chosen for BNN output. In the case of a Kendall and Gal kendall2017uncertainties style architecture, the model mean is directly parameterized as one of the model outputs. This fact may be exploited to avoid a sample average computation of the average. A sample variance calculation may then be performed on the collection of mean predictions. The Monte Carlo sampling and sample variance calculation is subject to the error prescribed by the Central Limit Theorem, again $\mathcal{O}(1/\sqrt{N_{e}})$
@@
3:     Compute BNN prediction of $\mathbf{y}_{\boldsymbol{\mu}}$ corresponding to $\mathbb{E}_{\mathbf{w_{j}}\sim q(\mathbf{w}|\theta)}[\mathbf{y}|\mathbf{x},\boldsymbol{D}]$ and append to collection $\{\mathbf{y}_{\boldsymbol{\mu}_{i}}\}^{N_{e}}_{i}$
@@
5:Compute the sample variance of collection $\{\mathbf{y}_{\boldsymbol{\mu}_{i}}\}^{N_{e}}_{i}$ corresponding to $\operatorname{Var}(\mathbb{E}_{q(\mathbf{w}|\theta)}[\mathbf{y}|\mathbf{x},\boldsymbol{D}])$
@@
3:     Compute BNN prediction of $\mathbf{y}_{\boldsymbol{\sigma}}$ corresponding to $\operatorname{Var}(\mathbf{y}|\mathbf{x},\boldsymbol{D})$ and append to collection $\{\mathbf{y}_{\boldsymbol{\sigma}_{i}}\}^{N_{a}}_{i}$
@@
5:Compute the mean of collection $\{\mathbf{y}_{\boldsymbol{\sigma}_{i}}\}^{N_{a}}_{i}$ corresponding to $\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\operatorname{Var}(\mathbf{y}|\mathbf{x},\boldsymbol{D})\right]$

### v2

@@ 3.3 Validation and Interpretation in Physical Space
The predicted total filtered progress variable dissipation rates, $\widetilde{\chi}=\chi_{\widetilde{C}}+\chi_{\rm SFS}$ and associated uncertainties (epistemic and aleatoric) are visualized for two different flames: an n-heptane flame with Karlovitz number $7$ [58] which was not included in the dataset (referred to as $B_{\rm Le}$), and a n-heptane flame with Karlovitz number $256$ which was included in the dataset (referred to as $D_{\rm Le}$). The prediction results and the ground truth data are shown for filter width 4 (Fig. 9) and 16 (Fig. 10).
@@ 3.4 Extrapolation Behavior
$\|\mu_{\rm OOD}-\mathbb{E}_{q_{c}(\mathbf{w}|\theta)}\left(\mathbb{E}\left[\mathbf{y}|\mathbf{x}\right]\right)\|_{2}/\sqrt{N},$ (11)
where the $L_{2}$ norm is computed over the set of synthetic data $\boldsymbol{D_{\rm OOD}}$ and $\mathbf{x}$ $N$ is the number of synthetic data points in $\boldsymbol{D_{\rm OOD}}$, and the normalization $1/\sqrt{N}$ accounts for the different synthetic dataset sizes. Likewise, the normalized $L_{2}$ error for the second moment is computed as
$\|\sigma_{\rm OOD}^{2}-\mathbb{E}_{q_{c}(\mathbf{w}|\theta)}\left(\operatorname{Var}\left[\mathbf{y}|\mathbf{x}\right]\right)\|_{2}/\sqrt{N}.$ (12)
@@
With the addition of increasing amounts of synthetic data, the model is tasked with representing a larger space and it is natural to expect that the performance should degrade on the original dataset. Figure 12 shows the effect of the amount of synthetic data on the relative $L_{2}$ error on the original dataset. For a given model $M$, the predictive mean is noted as $P_{M}=\mathbb{E}_{q_{M,c}(\mathbf{w}|\theta)}\left[\mathbf{y}|\mathbf{x}\right]$. The error metric considered is a relative error between the mean predictions of a model trained without synthetic data $P_{\rm ref}$ and a model trained with different amounts of synthetic data $P_{\rm OOD}$, as
@@
3.5 Out-of-Distribution Query Detection
The extrapolative capabilities enforced with the synthetic dataset can serve two purposes: first, they can ensure that extrapolation is at least as accurate as some low-fidelity model; second, they can serve as a marker for OOD queries. In this work, the synthetic data labeling was set arbitrarily rather than being guided by a low-fidelity model. Therefore the trained BNNs described in Sec. 3.4 best serve the second purpose: OOD query detection.
3.5.1 Detection Algorithm
Using a BNN trained with synthetic OOD data, the OOD query detection method can be done as described in Algo. 2.
Algorithm 2 Out-of-distribution query detection for input $\mathbf{x}$.
1:Define a distance threshold $T$
2:Predict $\chi_{\rm SFS,\mu}=\mathbb{E}_{q_{c}(\mathbf{w}|\theta)}\left(\mathbb{E}\left[\mathbf{y}|\mathbf{x}\right]\right)$ and $\chi_{\rm SFS,\sigma}=\operatorname{Var}_{q_{c}(\mathbf{w}|\theta)}\left(\mathbb{E}\left[\mathbf{y}|\mathbf{x}\right]\right)^{1/2}$
3:Compute the distance $d_{\rm OOD}=||(\chi_{\rm SFS,\mu},\chi_{\rm SFS,\sigma})-(\mu_{\rm OOD},\sigma_{\rm OOD})||_{2}$.
4:if $d_{\rm OOD}>T$ then
5:     $\mathbf{x}$ is in-distribution
6:else
7:     $\mathbf{x}$ is out-of-distribution
8:end if
In Algo. 2, the distance $d_{\rm OOD}$ to the arbitrary distribution of labels in the OOD region is used as a marker for whether or not an input $\mathbf{x}$ is OOD. Clearly the choice of the threshold $T$ impacts efficacy of the method. A systematic method to choose $T$ would depend on the labels adopted for the synthetic OOD data and is out-of-scope of this work. In the following, $T$ set to $0.6\times\sqrt{\mu_{\rm OOD}^{2}+\sigma_{\rm OOD}^{2}}$ which was chosen through a sensitivity analysis (See D).
3.5.2 Results
The OOD query detection is applied to the $D_{Le}$ and $B_{Le}$ flame contours already used in Sec. 3.3 for validation of the BNN predictions. The BNN model used is the one trained with $10^{7}$ synthetic OOD datapoints using the normalizing flow method (Algo. 1). Figure 13 (top) shows the in-distribution index that characterizes whether a query point (each point where the BNN is queried) is in or out-of-distribution. An in-distribution index is constructed with a value $1$ if $d_{\rm OOD}>T$ and 00 if $d_{\rm OOD}<T$. Figure 13 (bottom) shows the distance between the query point and its nearest neighbor in the training and testing datasets. If this distance (hereafter referred to as the the nearest neighbor metric) is large, then the query point is far from any data point seen by the BNN during training, and the query point should be considered to be OOD. The metric is used to assess whether the BNN accurately labels OOD query points.
(a) $B_{\rm Le}$ case.
(b) $D_{\rm Le}$ case.
Figure 13: Top: in-distribution index predicted by Algo. 2 for filterwidth of 4 and 16. The isocontour denotes the flame contour based on the progress variable value. Bottom: distance between the query point and the nearest neighbor of the query point in the training and testing dataset.
It is observed that the in-distribution index plots and the distance to nearest neighbor metric are in agreement. In particular, the BNN successfully labeled query points to be OOD when the distance to nearest neighbor is the largest. The OOD detection with the BNN can be done at the cost of BNN inference to compute the empirical averages of $\mathbb{E}\left[\mathbf{y}|\mathbf{x}\right]$ (Algo. 4) and $\mathbb{E}\left[\mathbf{y}|\mathbf{x}\right]$ (Algo. 5) and takes on the order of $\mathcal{O}(10^{-3})$ seconds on a single CPU. The computation of nearest neighbor metric required 108 CPUh (using Intel Xeon Gold Skylake CPUs). Therefore, the OOD query detection using a BNN is computationally advantageous compared with more naive methods.
Figure 13(a) (top) shows that the none of query points from the $B_{Le}$ flame case could be considered out-of-distribution. This is a surprising result given that the $B_{Le}$ flame was not included in the training dataset. Nevertheless, the phase-space spanned by the $B_{Le}$ flame is similar to the training and testing dataset. This result highlights that good predictive performances of data-based closure models on cases not included in the training dataset does not test the extrapolative abilities of data-based closure models. Instead, it may test whether the dataset was sufficiently large so that the data-based model can be used on reacting flows not included in the training dataset.
In turn, Fig. 13(b) (top) suggests that the case with filterwidth of 4 for the $D_{Le}$ flame contains OOD query points. This is, again, a surprising result given that the $D_{Le}$ flame was included in the training dataset. The nearest neighbor metric (Fig. 13(b), bottom) confirms that this finding is reasonable since the OOD query points are indeed far from any training or testing data points. This effect could be explained by the fact that the training data was preprocessed with an approximate uniform-in-phase-space sampling to eliminate redundant datapoints. However the method used in Ref. [46] has been found to be inaccurate since it tends to remove too many rare datapoints [25]. Fortunately, the OOD query is more prominent for the filterwidth 4 which is also where the contribution of $\chi_{\rm SFS}$ is negligible compared to $\chi_{\widetilde{C}}$. This explains why the OOD queries did not affect the reconstruction of $\widetilde{\chi}$ in this work (Sec. 3.3) or in Ref. [46].
4 Conclusions
@@
A strategy for enforcing a specific OOD behavior was proposed using synthetic data generation. It was found that using a uniform in-phase space data generation led to the best performances and that the performance in-distribution did not degrade until the synthetic dataset size approached that of the original dataset. The OOD behavior can be used as a marker that efficiently tests whether the inputs passed to the BNN are far or not from the data-distribution. It was found that the strategy proposed is computationally efficient for identifying OOD queries.
@@
5 CRediT Authorship Contribution Statement
GP: Methodology, Investigation, Software, Writing - original draft. MH: Conceptualization, Methodology, Investigation, Software, Supervision, Visualization, Writing - review & editing. SY: Conceptualization, Data curation, Funding acquisition, Supervision, Writing - review & editing.
@@ 6 Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0021110. This work was authored in part by the National Renewable Energy Laboratory (NREL), operated by Alliance for Sustainable Energy, LLC, for the U.S. Department of Energy (DOE) under Contract No. DE-AC36-08GO28308. This work was supported as part of DEGREES, an Energy Earthshot Research Center (EERC) funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences. SY was supported by the U.S. Department of Energy Office of Energy Efficiency and Renewable Energy Vehicle Technologies Office (VTO). The research was performed using computational resources sponsored by the Department of Energy’s Office of Energy Efficiency and Renewable Energy and located at the National Renewable Energy Laboratory. The views expressed in the article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes. This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
@@ Appendix B Computation of Uncertainties
In this appendix, algorithms for the computation of a predictive distribution (Alg. 3), the epistemic uncertainty (Alg. 4) and the aleatoric uncertainty (Alg. 5) are presented. Overall, the BNN allows for fast evaluation of the samples of weights, but appropriate averaging is needed to differentiate between epistemic and aleatoric uncertainties.
@@
Computation of the epistemic uncertainty relies on the parameterization chosen for BNN output. In the case of a Kendall and Gal [66] style architecture, the model mean is directly parameterized as one of the model outputs. This fact may be exploited to avoid a sample average computation of the average. A sample variance calculation may then be performed on the collection of mean predictions. The Monte Carlo sampling and sample variance calculation is subject to the error prescribed by the Central Limit Theorem, again $\mathcal{O}(1/\sqrt{N_{e}})$
@@
3:     Compute BNN prediction of $\mathbf{y}_{\boldsymbol{\mu}}$ corresponding to $\mathbb{E}[\mathbf{y}|\mathbf{x}]$ for the weights $\mathbf{w}_{j}$ and append to collection $\{\mathbf{y}_{\boldsymbol{\mu}_{i}}\}^{N_{e}}_{i}$
@@
5:Compute the sample variance of collection $\{\mathbf{y}_{\boldsymbol{\mu}_{i}}\}^{N_{e}}_{i}$ corresponding to $\operatorname{Var}_{q(\mathbf{w}|\theta)}(\mathbb{E}[\mathbf{y}|\mathbf{x}])$
@@
3:     Compute BNN prediction of $\mathbf{y}_{\boldsymbol{\sigma}}$ corresponding to $\operatorname{Var}(\mathbf{y}|\mathbf{x})$ for the weights $\mathbf{w}_{j}$ and append to collection $\{\mathbf{y}_{\boldsymbol{\sigma}_{i}}\}^{N_{a}}_{i}$
@@
5:Compute the mean of collection $\{\mathbf{y}_{\boldsymbol{\sigma}_{i}}\}^{N_{a}}_{i}$ corresponding to $\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\operatorname{Var}(\mathbf{y}|\mathbf{x})\right]$
@@ Appendix C Effect of an Isotropic Gaussian Prior on Prediction Quality
Appendix D Effect of distance threshold on out of distribution query detection
In Algo. 2, a key variable for the OOD of query points is the threshold $T$ which discriminates between query points in the OOD regime and in-distribution regime. The value of $T$ depends on the magnitude of the distance metric $d_{\rm OOD}$, which itself depends on the magnitude of $\mu_{\rm OOD}$ and $\sigma_{\rm OOD}$. In Sec. 3.5, the value of $T$ was therefore defined as $\alpha\sqrt{\mu_{\rm OOD}^{2}+\sigma_{\rm OOD}^{2}}$, where it was chosen $\alpha=0.6$. To define a reasonable value of $\alpha$, a sensitivity analysis was performed: the value of $\alpha$ was varied until the fields of in-distribution index became insensitive to $\alpha$.
(a) $B_{\rm Le}$ case.
(b) $D_{\rm Le}$ case.
Figure 16: In-distribution index predicted by Algo. 2 for filterwidth of 4 and 16. Top: $\alpha=0.4$. Middle: $\alpha=0.6$. Bottom: $\alpha=0.8$. The isocontour denotes the flame contour based on the progress variable value.
Figure 16 shows that the value of $\alpha$ affects the in-distribution index fields. However, the in-distribution index fields is significantly less sensitive to $\alpha$ if $\alpha<0.6$ than if $\alpha>0.6$, which explains the choice adopted for $\alpha$ in Sec. 3.5.